# The Alexander polynomial of twisted torus knots

### v1

@@ 1. Introduction
Given a quadruple of integers $(p,q;r,s)$ such that $\gcd(p,q)=1$, $2\leq r<p$ and $s\in\mathbb{Z}$, the twisted torus knot $T(p,q;r,s)$ is a generalization of torus knots, achieved by applying $s$ full twists to $r$ adjacent strands of the standard diagram of the torus knot $T(p,q)$. This knot can also be described as the closure of the braid corresponding to the braid word
@@
with $p$ strands. Figure 1 illustrates an example of the twisted torus knot $T(7,4;3,2)$. Note that $T(p,q;r,s)$ is isotopic to $T(q,p;r,s)$, and $T(p,-q;r,s)$ is isotopic to the mirror image of $T(p,q;r,-s)$. Considering our main interests in this paper, such as the Alexander polynomial and the genus of a knot, are invariant under the mirror imaging of a knot, we will assume that $p>q>0$ with $\gcd(p,q)=1$, $1<r<p$, and $s\in\mathbb{Z}$ for twisted torus knot $T(p,q;r,s)$ throughout this paper.
@@
However, not much is known about the polynomial knot invariants for twisted torus knots. Since some twisted torus knots are known to be the unknot, or more generally, torus knots as mentioned above, the Alexander polynomials of twisted torus knots in those families are implicitly known. In [Mor06], Morton provided an explicit formula for the Alexander polynomial of twisted torus knots of the type $T(p,q;2,s)$. His main technique involved the Burau representation alongside an inductive argument based on the skein relation of the Alexander polynomial. If $r=2$, the other skein pairs or a twisted torus knot are also twisted torus knots or links, allowing the use of a recursive formula to obtain the Alexander polynomial of the twisted knot. However, this technique does not work for twisted torus knots with $r>2$, as the other pairs in the skein relation are no longer twisted torus knots or links. The knot Floer homology, which refines Alexander polynomial invariant, for certain family of twisted torus knots is studied in [Vaf15]. In [BD23], the Jones polynomial of twisted torus knots of the type $T(p,q;2,s)$ is computed.
@@
$\tilde{X}(t)$ $=1-(1-t^{rs})\sum_{i=1}^{m}t^{\overline{k_{i}}p+(i-1)rs}-t^{\overline{k^{\prime}}p+mrs},$
$Y(t)$ $=1-(1-t^{rs})\sum_{i=1}^{r-1}t^{Q_{i}q+(i-1)rs}-t^{pq+(r-1)rs},\text{ and}$
$\tilde{Y}(t)$ $=1-(1-t^{rs})\sum_{i=1}^{m}t^{Q_{i}q+(i-1)rs}-t^{Q^{\prime}q+mrs}.$
For cases where the upper bound of the summation $m$ equals zero, we assume that the summation term is zero.
@@
One interesting topic in low-dimensional topology is the classification of all knots in $S^{3}$ that produce lens spaces through integral surgery long the knots, known as the Berge conjecture. In [Mor06], Morton showed the existence of an infinite family of twisted torus knots of the type $T(p,q;2,s)$ whose Alexander polynomials have coefficients that are neither ±$\pm 1$ nor 00. Consequently, by the result of Ozsváth and Szabó’s [OS05, Corollary 1.3], these knots do not admit lens space surgeries; in other words, any integral surgery on these knots does not produce lens spaces. Using our formula, we have also identified an infinite family of twisted torus knots exhibiting the same property.
@@
If $n-2\geq r\geq 2$ and $s\geq 2$, then the Alexander polynomial of the twisted torus knot $T(2n+1,n,r,s)$ has a coefficient whose absolute value is $\geq 2$. Therefore, $T(2n+1,n,r,s)$ does not admit lens space surgery (more generally, this knot is not an $L$-space knot).
@@
It is interesting to explore how the Alexander polynomial serves as a powerful invariant for twisted torus knots. It is not difficult to observe that there are many pairs of twisted torus knots with different parameters that are isotopic, even in the case when $s>0$. For example, $T(1+3n,3;2,1+m)$ and $T(3+2m,2;3,n)$ are isotopic for each $n>0$ and $m>0$. We implemented code to compute the Alexander polynomial of twisted torus knots using our formula. Specifically, let $\mathcal{T}$ denote the set of $7,450$ twisted torus knots with $3\leq p\leq 20$, $0<q<p$ with $\gcd(p,q)=1$, $1<r<p$ and $1\leq s\leq 5$. We found that each pair in $\mathcal{T}$ with the same Alexander polynomial also has the same volume (up to an error $<10^{-8}$ with SnapPy’s results) if they are hyperbolic knots (determined by SnapPy [CDGW]). Conversely, any pair of two hyperbolic twisted torus knots in $\mathcal{T}$ with the identical volume (determined by SnapPy) also have the same Alexander polynomial. These observations lead us to pose the following question.
@@
Is Alexander polynomial a complete invariant for (positively) twisted torus knots?
@@
Twisted torus knots can be generalized or may overlap with larger families of knots and links, such as twisted torus links, links obtained by concatenating two braid words for torus links, and $T$-links (equivalently Lorenz links), which are links formed by concatenating multiple braid words for torus links $T(p_{i},q_{i})$ with $q_{i}>0$ [BK09]. We anticipate that the techniques used in this work can also be extended to these families to compute their Alexander polynomials.
@@
Remark.
Twisted torus knots are originally defined for the parameter $2\leq r\leq p+q$ (see [Dea03, Chapter 3.1]. However, in this article, we restrict our attention to the case where $2\leq r<\max\{p,q\}$.
@@ Acknowledgements
The authors would like to thank Thiago de Paiva Souza for his interest in this work and for helpful discussions. Kyungbae Park was supported by 2020 Research Grant from Kangwon National University and the National Research Foundation of Korea (NRF) grant funded by the Korea government (No.2022R1F1A1071673).
@@ 2. Fundamental Group of Twisted Torus Knots
$k_{i}:=\left|\{l\in R|Q_{i-1}\leq l<Q_{i}\}\right|\text{ for }i=1,\dots,r-1,\text{ and }$
@@
$\overline{k_{i}}:=k_{1}+\dots+k_{i}\text{ for }i=1\dots r-1,\quad\text{and}$
@@
Let $p$ and $q$ are relatively prime with $0<q<p$, and $2\leq r<p$ and $s\in\mathbb{Z}$. Then the knot group of the twisted torus knot $T(p,q;r,s)$ has the following presentation.
@@
Wheres, $b$ intersects $\alpha_{2}$ once during its travel, and the orientation of this intersection is opposite to that of $\beta_{2}$. Hence,
@@
$\xi_{n}=\begin{cases}xy^{s}&\text{if }t_{n}\leq q\text{ and }u_{n}<r\\ y^{s}&\text{if }t_{n}>q\text{ and }u_{n}<r\\ x&\text{if }t_{n}\leq q\text{ and }u_{n}\geq r\\ 1&\text{if }t_{n}>q\text{ and }u_{n}\geq r\\ \end{cases}$
@@
$\xi_{n}=\begin{cases}z^{-1}w^{-1}&\text{if}\quad u_{n}<r\\ z^{-1}&\text{if}\quad u_{n}\geq r\\ \end{cases}$
@@
The same argument applies to obtaining $i(a)$ and $j(a)$. However, in this case, the loop $a$ meets $\partial A$ exactly $m:=|\{[nq^{-1}|n=0,\dots,r-1,[nq^{-1}]<[rq^{-1}]\}|$ times, since $a$ returns to the base point when it passes through the sector of $L$ labeled $r$. Note that $Q^{\prime}:=[rq^{-1}]$ is the solution $n$ of the equation $u_{n}=r$. Considering this situation, we have
$i(a)=\prod_{n=1}^{m}\lambda_{n}\cdot y$
where
$\lambda_{n}=\begin{cases}xy^{s}&\text{if }t_{n}>p-q\text{ and }u_{n}<r\\ y^{s}&\text{if }t_{n}\leq p-q\text{ and }u_{n}<r\\ x&\text{if }t_{n}>p-q\text{ and }u_{n}\geq r\\ 1&\text{if }t_{n}\leq p-q\text{ and }u_{n}\geq r\\ \end{cases}$
for $n=1,\dots,m$, and hence
@@
$j(a)=\prod_{n=1}^{m}\tau_{n},$
where
$\tau_{n}=\begin{cases}z^{-1}w^{-1}&\text{if}\quad u_{n}<r\\ z^{-1}&\text{if}\quad u_{n}\geq r.\\ \end{cases}$
@@ 3.2. Alexander polynomial of twisted torus knots
We also carry all notation introduced in Section 2. Before proceeding further, note the following identity, which will be useful later.
@@
Recall that $Q^{\prime}=[rq^{-1}]$ and $\overline{k^{\prime}}$ denotes the number of elements in
$\{[-nq^{-1}]|n=1,\dots,q,\,[-nq^{-1}]<Q^{\prime}\}.$
Observe that $Q^{\prime}$ is the minimal positive integer $m$ such that $mq\equiv r$ modulo $p$. Hence, there exists a positive integer $t$ such that $Q^{\prime}q=tp+r$. We claim that $t=\overline{k^{\prime}}$. For each $1\leq v\leq t$, we have $m_{v}q$, which is the greatest multiple of $q$ less than $vp$. Therefore, since $1\leq vp-m_{v}q\leq q$, $t$ equals to the number of solutions of $x$ such that $0<x<Q^{\prime}$ and $x=[-nq^{-1}]$ for some $1\leq n\leq q$, which corresponds to $\overline{k^{\prime}}$. ∎
@@
where we applied Lemmma 9 for the third identity. Similarly, we have
@@
Then, the minors $m_{1}$, $m_{2}$, $m_{3}$ and $m_{4}$ of the above will be the following
@@
It is straightforward to observe that $t^{rs}(1-t^{r})m_{4}=(1-t^{rs})m_{2}$. We further claim that similar identities exists for $m_{1}$ and $m_{3}$, each relating to $m_{2}$.
@@ 4. Degree of the Alexander polynomial of Twisted Torus Knots
$A=$ $\ t^{\overline{k^{\prime}}p+mrs}+(1-t^{rs})\sum_{i=m+1}^{r-1}t^{\overline{k^{\prime}}p+(i-1)rs}-t^{pq+(r-1)rs},\text{ and}$
@@
(7) $\tilde{X}Y-X\tilde{Y}=\tilde{X}B-A\tilde{Y}.$
@@
Given that $|rs|<q<p$, it follows that $|(i-1)rs|<Q_{i}q$ and $|(i-1)rs|<{k_{i}}p$ for all $i$. Consequently, the highest and the lowest powers of $t$ in $\tilde{X}Y-X\tilde{Y}$ remain unchanged from the case when $s>0$. ∎
@@
Suppose $s<0$, and $r|s|>d_{i}q$ and $r|s|>\overline{k_{i}}p$ for all $i=1,\dots,r-1$. Then
@@
    $\deg(\Delta_{T(p,q;r,s)})=(k_{1}-\overline{k_{r-1}}-1)p+(Q^{\prime}-1)q-(r-1)(rs+1)$ if $m=0$.
@@
    $\deg(\Delta_{T(p,q;r,s)})=(p-1)(q-1)+(Q_{1}-Q^{\prime}-Q_{r-1})q-(r-1)rs$ if $m=r-1$.
@@
    $\deg(\Delta_{T(p,q;r,s)})=(\overline{k_{m+1}}-\overline{k_{r-1}}-1)p+(Q_{1}-Q_{m}-1)q-(r-1)(rs+1)$ if $10<m<r-$ 1.
@@
If $m=0$, then $\tilde{X}$ and $\tilde{Y}$ are $1-t^{\overline{k^{\prime}}p}$ and $1-t^{Q^{\prime}q}$ respectively. Since $|rs|>Q_{i}q$, $|rs|>\overline{k_{i}}p$ for all $i$, then the highest degrees in $\tilde{X}Y$ and $X\tilde{Y}$ are $\overline{k^{\prime}}p+Q_{1}q$ and $Q^{\prime}q+k_{1}p$ respectively. By Lemma 9 the highest degree in $\tilde{X}Y-X\tilde{Y}$ is $Q^{\prime}q+k_{1}p$. While the lowest degrees in $\tilde{X}Y$ and $X\tilde{Y}$ are $Q_{r-1}q+(r-2)rs+rs$ and $\overline{k_{r-1}}p+(r-2)rs+rs$ respectively. Since $Q_{i}q>\overline{k_{i}}p$ for all $i$ then the lowest degree in $\tilde{X}Y-X\tilde{Y}$ is $\overline{k_{r-1}}p+(r-2)rs+rs$. Hence,
$\text{deg}(\Delta_{T(p,q;r,s)})=(k_{1}-\overline{k_{r-1}}-1)p+(Q^{\prime}-1)q-(r-1)(rs+1).$
If $m=r-1$ by Theorem 1 we have:
$\tilde{X}$ $=1-(1-t^{rs})\sum_{i=1}^{r-1}t^{\overline{k_{i}}p+(i-1)rs}-t^{\overline{k^{\prime}}p+(r-1)rs},$
$\tilde{Y}$ $=1-(1-t^{rs})\sum_{i=1}^{r-1}t^{Q_{i}q+(i-1)rs}-t^{Q^{\prime}q+(r-1)rs}$
Then $X=\tilde{X}-A$ and $Y=\tilde{Y}-B$ , where
$A=t^{pq+(r-1)rs}-t^{\overline{k^{\prime}}p+(r-1)rs}$

### v3

@@ 1. Introduction
Given a quadruple of integers $(p,q;r,s)$ such that $\gcd(p,q)=1$, $1<r\leq|p|+|q|$ and $s\in\mathbb{Z}$, the twisted torus knot $T(p,q;r,s)$ is a generalization of torus knots, achieved by applying $s$ full twists to $r$ adjacent strands in the standard diagram of torus knot $T(p,q)$ on the flat torus. In particular, when $r<p$, which is the case we mainly consider in this paper, this knot can also be described as the closure of the braid corresponding to the braid word
@@
with $p$ strands. Figure 1 illustrates an example of the twisted torus knot $T(7,4;3,2)$. Note that $T(p,q;r,s)$ with $r\leq q$ is isotopic to $T(q,p;r,s)$, and $T(p,-q;r,s)$ is isotopic to the mirror image of $T(p,q;r,-s)$. Considering our main interests in this paper, such as the Alexander polynomial and the genus of a knot, are invariant under the mirror imaging of a knot, we will assume that $p>q>0$ with $\gcd(p,q)=1$, $1<r<p$, and $s\in\mathbb{Z}$ for twisted torus knot $T(p,q;r,s)$ throughout this paper.
@@
However, not much is known about the polynomial knot invariants for twisted torus knots. Since some twisted torus knots are known to be the unknot, or more generally, torus knots as mentioned above, the Alexander polynomials of twisted torus knots in those families are implicitly known. In [Mor06], Morton provided an explicit formula for the Alexander polynomial of twisted torus knots of the type $T(p,q;2,s)$. His main technique involved the Burau representation alongside an inductive argument based on the skein relation of the Alexander polynomial. If $r=2$, the other skein pairs of a twisted torus knot are also twisted torus knots or links, allowing the use of a recursive formula to obtain the Alexander polynomial of the twisted knot. However, this technique does not work for twisted torus knots with $r>2$, as the other pairs in the skein relation are no longer twisted torus knots or links. The knot Floer homology, which refines Alexander polynomial invariant, for twisted torus knots with $p=qk\pm 1$ is studied in [Vaf15]. In [BD24], the Jones polynomial of twisted torus knots of the type $T(p,q;2,s)$ is computed.
@@
$\tilde{X}(t)$ =$=\begin{cases}1-t^{\overline{k^{\prime}}p},&\text{if }m=0,\\ 1-(1-t^{rs})\displaystyle\sum_{i=1}^{m}t^{\overline{k_{i}}p+(i-1)rs}-t^{\overline{k^{\prime}}p+mrs},&\text{otherwise,}\end{cases}$
$Y(t)$ $=1-(1-t^{rs})\sum_{i=1}^{r-1}t^{Q_{i}q+(i-1)rs}-t^{pq+(r-1)rs},$
$\tilde{Y}(t)$ =$=\begin{cases}1-t^{Q^{\prime}q},&\text{if }m=0,\\ 1-(1-t^{rs})\displaystyle\sum_{i=1}^{m}t^{Q_{i}q+(i-1)rs}-t^{Q^{\prime}q+mrs},&\text{otherwise.}\end{cases}$
@@
One interesting topic in low-dimensional topology is the classification of all knots in $S^{3}$ that produce lens spaces through integral surgery along the knots, known as the Berge conjecture. In [Mor06], Morton showed the existence of an infinite family of twisted torus knots of the type $T(p,q;2,s)$ whose Alexander polynomials have coefficients that are neither ±$\pm 1$ nor 00. Consequently, by the result of Ozsváth and Szabó’s [OS05, Corollary 1.3], these knots do not admit lens space surgeries; in other words, any integral surgery on these knots does not produce lens spaces. In [Vaf15], Vafaee classified twisted torus knots of the type $T(qk\pm 1,q;r,s)$ that admit $L$-space surgery, a generalization of lens space surgery in the context of Heegaard Floer theory. Using our formula, one can identify infinite families of twisted torus knots that do not admit lens space surgery through a simple computation. The following is one such new example.
@@
For each $n\geq 0$ and $s\geq 2$, the Alexander polynomial of the twisted torus knot $T(9+2n,7+2n;3,s)$ has a coefficient with an absolute value of at least $2$. Consequently, this knot does not admit lens space surgery (and, more generally, is not an $L$-space knot).
@@
It is interesting to explore how the Alexander polynomial serves as a powerful invariant for twisted torus knots. It is not difficult to observe that there are many pairs of twisted torus knots with different parameters that are isotopic, even in the case when $s>0$. For example, $T(1+3n,3;2,1+m)$ and $T(3+2m,2;3,n)$ are isotopic for each $n>0$ and $m>0$. We implemented Python code to compute the Alexander polynomial of twisted torus knots using our formula (see the second author’s homepage). Specifically, let $\mathcal{T}$ denote the set of $7,450$ twisted torus knots with $3\leq p\leq 20$, $0<q<p$ with $\gcd(p,q)=1$, $1<r<p$ and $1\leq s\leq 5$. We found that each pair in $\mathcal{T}$ with the same Alexander polynomial also has the same volume (up to an error $<10^{-8}$ with SnapPy’s results) if they are hyperbolic knots (determined by SnapPy [CDGW]). Conversely, any pair of two hyperbolic twisted torus knots in $\mathcal{T}$ with the identical volume (determined by SnapPy) also has the same Alexander polynomial. These observations lead us to pose the following question.
@@
Is Alexander polynomial a complete invariant for positively twisted torus knots?
Here “positively twisted” refers to the case when $s>0$. The question above is not expected to hold for negatively twisted torus knots, as we identified pairs of negatively twisted torus knots that share the same Alexander polynomial but have different volumes according to SnapPy. For example, $T(4,3;2,-5)$ and $T(6,5;4,-2)$ form such a pair.
@@
Twisted torus knots are originally defined for the parameter $2\leq r\leq|p|+|q|$ (see [Dea03, Chapter 3.1]). Although we restrict our attention to the case where $2\leq r<\max\{|p|,|q|\}$, similar arguments can be applied to the general case, though they would yield a different type of formula.
Twisted torus knots can also be generalized or may overlap with larger families of knots and links, such as twisted torus links, links obtained by concatenating two braid words for torus links, and $T$-links (equivalently Lorenz links), which are links formed by concatenating multiple braid words for torus links $T(p_{i},q_{i})$ with $q_{i}>0$ [BK09]. We anticipate that the techniques used in this work can also be extended to these families to compute their Alexander polynomials.
@@ Acknowledgements
The authors thank the referee for valuable suggestions that improved the manuscript. They also thank Thiago de Paiva for his interest and helpful discussions. Kyungbae Park was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (No. RS-2022-NR073368).
@@ 2. Fundamental Group of Twisted Torus Knots
$k_{i}:=\left|\{l\in R|Q_{i-1}\leq l<Q_{i}\}\right|\text{ for }i=1,\dots,r-1,$
and
@@
$\overline{k_{i}}:=k_{1}+\dots+k_{i}\text{ for }i=1\dots r-1,$
and
@@
Example.
We give an example illustrating the computation of the above numbers for an infinite family of triples $(p,q,r)$, which will be used in Section 5. Let $(p,q,r)=(9+2\ell,7+2\ell,3)$ where $\ell$ is a nonnegative integer. Note that $[q^{-1}]=[(7+2\ell)^{-1}]=4+\ell$. Therefore
$Q=\left\{\left[n(4+\ell)\right]|n=0,1,2\right\}=\{0,4+\ell,8+2\ell\},$
so we have
$Q_{0}=0,\quad Q_{1}=4+\ell,\quad\text{and}\quad Q_{2}=2(4+\ell),$
and
$d_{1}=d_{2}=4+\ell\quad\text{and}\quad d_{3}=p-Q_{2}=1.$
We also have
$R=\left\{[-n(4+\ell)]|n=1,2,\dots,(7+2\ell)\right\}=\{1,2,\dots,3+\ell,5+\ell,6+\ell\dots,8+2\ell\},$
and hence
$k_{1}=k_{2}=3+\ell,\quad\text{and}\quad k_{3}=1.$
Thus
$\overline{k_{1}}=3+\ell\quad\text{and}\quad\overline{k_{2}}=2(3+\ell).$
Since $Q^{\prime}=[3(4+\ell)]=3+\ell$ and so $Q_{0}<Q^{\prime}<Q_{1}$, we obtain
$m=0\quad\text{and}\quad\overline{k^{\prime}}=2+\ell.$
@@
Let $p$ and $q$ be relatively prime with $0<q<p$, and $2\leq r<p$ and $s\in\mathbb{Z}$. Then the knot group of the twisted torus knot $T(p,q;r,s)$ has the following presentation.
@@
Whereas, $b$ intersects $\alpha_{2}$ once during its travel, and the orientation of this intersection is opposite to that of $\beta_{2}$. Hence,
@@
$\xi_{n}=\begin{cases}xy^{s}&\text{if }t_{n}<q\text{ and }u_{n}<r\\ y^{s}&\text{if }t_{n}\geq q\text{ and }u_{n}<r\\ x&\text{if }t_{n}<q\text{ and }u_{n}\geq r\\ 1&\text{if }t_{n}\geq q\text{ and }u_{n}\geq r\\ \end{cases}$
@@
$\eta_{n}=\begin{cases}z^{-1}w^{-1}&\text{if}\quad u_{n}<r\\ z^{-1}&\text{if}\quad u_{n}\geq r\\ \end{cases}$
@@
The same argument applies to obtaining $i(a)$ and $j(a)$. However, in this case, the loop $a$ meets $\partial A$ exactly $Q^{\prime}$ times since $Q^{\prime}:=[rq^{-1}]$ is the solution $n$ of the equation $u_{n}=r$ and $a$ returns to the base point when it passes through the sector of $L$ labeled $r$. Considering this situation, we have
$i(a)=\prod_{n=1}^{Q^{\prime}}\xi_{n}\cdot y.$
Hence,
@@
$j(a)=\prod_{n=1}^{Q^{\prime}}\eta_{n}.$
@@ 3.2. Alexander polynomial of twisted torus knots
We also carry all notation introduced in Section 2. Before proceeding further, note the following identities, which will be useful later.
@@
For each $i=0,\dots,r-1$, if $Q_{i}=[nq^{-1}]$ for some $n\in\{0,\dots,r-1\}$, then
$Q_{i}q=\overline{k_{i}}p+n,$
and
@@
We provide a proof of the last identity, as the others follow in a similar manner. Recall that $Q^{\prime}=[rq^{-1}]$ and $\overline{k^{\prime}}$ denote the number of elements in
$\{[-\ell q^{-1}]|\ell=1,\dots,q,\,[-\ell q^{-1}]<Q^{\prime}\}.$
Observe that $Q^{\prime}$ is the minimal positive integer $m$ such that $mq\equiv r$ modulo $p$. Hence, there exists a positive integer $t$ such that $Q^{\prime}q=tp+r$. We claim that $t=\overline{k^{\prime}}$. For each $1\leq v\leq t$, we have $m_{v}q$, which is the greatest multiple of $q$ less than $vp$. Therefore, since $1\leq vp-m_{v}q\leq q$, $t$ equals the number of solutions of $x$ such that $0<x<Q^{\prime}$ and $x=[-\ell q^{-1}]$ for some $1\leq\ell\leq q$, which corresponds to $\overline{k^{\prime}}$.
@@
where we applied Lemma 9 for the third identity. Similarly, we have
@@
Then, the minors $m_{1}$, $m_{2}$, $m_{3}$ and $m_{4}$ of the above will be the following:
@@
It is straightforward to observe that $t^{rs}(1-t^{r})m_{4}=(1-t^{rs})m_{2}$. We further claim that similar identities exist for $m_{1}$ and $m_{3}$, each relating to $m_{2}$.
@@ 4. Degree of the Alexander polynomial of Twisted Torus Knots
$A=\ t^{\overline{k^{\prime}}p+mrs}+(1-t^{rs})\sum_{i=m+1}^{r-1}t^{\overline{k_{i}}p+(i-1)rs}-t^{pq+(r-1)rs},$
and
@@
(7) $\tilde{X}Y-X\tilde{Y}=\tilde{X}B-\tilde{Y}A.$
@@
Given that $|rs|<q<p$, it follows that $|rs|<d_{i}q$ and $|rs|<{k_{i}}p$ for all $i$. Consequently, the highest and the lowest powers of $t$ in $\tilde{X}Y-X\tilde{Y}$ remain unchanged from the case when $s>0$. ∎
@@
Suppose $s<0$, and $|rs|>d_{i}q$ and $|rs|>k_{i}p$ for all $i=1,\dots,r-1$. Then
@@
    $\deg\left(\Delta_{T(p,q;r,s)}\right)=(k_{1}-\overline{k_{r-1}}-1)p+(Q^{\prime}-1)q-(r-1)(rs+1)$ if $m=0$.
@@
    $\deg\left(\Delta_{T(p,q;r,s)}\right)=(p-1)(q-1)+(Q_{1}-Q^{\prime}-Q_{r-1})q-(r-1)rs$ if $m=r-1$.
@@
    $\deg\left(\Delta_{T(p,q;r,s)}\right)=(\overline{k_{m+1}}-\overline{k_{r-1}}-1)p+(Q_{1}-Q_{m}-1)q-(r-1)(rs+1)$ if $10<m<r-$ 1.
@@
If $m=0$, then $\tilde{X}=1-t^{\overline{k^{\prime}}p}$ and $\tilde{Y}=1-t^{Q^{\prime}q}$. Since $|rs|>d_{i}q$ and $|rs|>k_{i}p$ for all $i$, the highest powers of $t$ in $\tilde{X}Y$ and $X\tilde{Y}$ are $Q_{1}q+\overline{k^{\prime}}p$ and $Q^{\prime}q+k_{1}p$ respectively. By Lemma 9 the highest power in $\tilde{X}Y-X\tilde{Y}$ is $Q^{\prime}q+k_{1}p$. Similarly the lowest powers in $\tilde{X}Y$ and $X\tilde{Y}$ are $Q_{r-1}q+(r-1)rs$ and $\overline{k_{r-1}}p+(r-1)rs$ respectively. Again by Lemma 9, the lowest power in $\tilde{X}Y-X\tilde{Y}$ is $\overline{k_{r-1}}p+(r-1)rs$. Hence, considering the degree of the formula in Theorem 1,
$\text{deg}\left(\Delta_{T(p,q;r,s)}\right)=(k_{1}-\overline{k_{r-1}}-1)p+(Q^{\prime}-1)q-(r-1)(rs+1).$
If $m=r-1$, we have:
$\tilde{X}=1-(1-t^{rs})\sum_{i=1}^{r-1}t^{\overline{k_{i}}p+(i-1)rs}-t^{\overline{k^{\prime}}p+(r-1)rs}$